What is the standard form of homogeneous differential equation of order n.
1
Expert's answer
2020-06-17T18:57:07-0400

HOMOGENEOUS DIFFERENTIAL EQUATION

An equation of the form

xn"\frac {d^ny}{dx^n}" + P1xn-1"\frac {d^{n-1}y}{dx^{n-1}}" + P2xn-2 "\frac {d^{n-2}y}{dx^{n-2}}" + ............+ Pn-1x"\frac {dy}{dx}" +Pny= X, where P1, P2, P3, ......,Pn are constants and X is a function of x alone , is homogeneous differential equation of order n .

This is the standard form of homogeneous differential equation of order n .